OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? the line perpendicular to y = 1/3x + 5 through (2, 1) A. y = 1/3x + 7 B. y = -1/3x + 7 C. y = -3x + 7 D. y = 3x + 7

OpenStudy (anonymous):

so lets find the slope perpendicular to this equation first

OpenStudy (anonymous):

dontknownomth is this the same process?

OpenStudy (anonymous):

or do i need two y's and two x's

OpenStudy (anonymous):

the last problem you had two points so this process is different

OpenStudy (anonymous):

so whats your slope in this equation?

OpenStudy (anonymous):

y=1/3x=5 i believe i could type that into desmos and see the pints it goes through?

OpenStudy (anonymous):

points*

OpenStudy (anonymous):

well now your looking for the perpendicular line

OpenStudy (anonymous):

i meant to put y=1/3x+5 btw and what process would i take to find this out?

OpenStudy (anonymous):

Here's a clear example

OpenStudy (anonymous):

y = 1/3x + 5 through (2, 1) STEP #1--So what is the your slope?

OpenStudy (anonymous):

that made me more confused

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

okay STEP #2- Find your perpendicular slope Do you know how to find it?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

To find your perpendicular slope is to "flip" your original slope and change the sign

OpenStudy (anonymous):

so 1/3 flipped is?

OpenStudy (anonymous):

3/1

OpenStudy (anonymous):

Yes!! and no change the sign to have your perpendicular slope

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

change the sign?

OpenStudy (anonymous):

so its + 3 but if we change the sign what will it be?

OpenStudy (anonymous):

-3/1

OpenStudy (anonymous):

well -3

OpenStudy (anonymous):

Yes! so our perpendicular slope is -3 to the original slope 1/3

OpenStudy (anonymous):

so to find the perpendicular slope we "flip" and change the sign

OpenStudy (anonymous):

which we just did

OpenStudy (anonymous):

right

OpenStudy (anonymous):

okay now comes STEP #3 We plug our new slope into the point slope formula Do you know what the point slope formula is?

OpenStudy (anonymous):

no clue

OpenStudy (anonymous):

Point Slope Formula: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

so whats our \[(x _{1},y _{1}) \] & m

OpenStudy (anonymous):

y-1=-3(x-2)

OpenStudy (anonymous):

Correct!!! now solve for y

OpenStudy (anonymous):

would i just be left with -3?

OpenStudy (anonymous):

where does the +5 come in

OpenStudy (anonymous):

nope okay i'm going to help you solve so we want what variable alone?

OpenStudy (anonymous):

the +5 has nothing to do with this perpendicular equation even though its part of the original equation

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so you found this equation so we want y alone so what do we have to do first with the -3? y-1=-3(x-2)

OpenStudy (anonymous):

move beside (x-2)

OpenStudy (anonymous):

you mean distribute to (x-2)

OpenStudy (anonymous):

distribute meaning?

OpenStudy (anonymous):

-3-2?

OpenStudy (anonymous):

(-3-2)

OpenStudy (anonymous):

Distributing=multiplying the number exactly outside the parenthesis into the parenthesis